OpenStudy (anonymous):

Shelby is baking muffins the require 1/2 cup vegetable oil. She only had 1/3 cup of vegetable oil left. How much more dose she need?

OpenStudy (ja1):

Ok so we need the LCD. can you find that for me?

OpenStudy (anonymous):

the answer is 1/6 because 3 and 2 both go into 6th's 1/3= 2/6 1/2=3/6 3/6-2/6=1/6